OpenStudy (anonymous):

19 - (2x + 3) = 2(x + 3) + x.....the anwser i got was -10. is that not the right anwser?

OpenStudy (anonymous):

That is not the correct answer. x=2 is the answer First distribute the negative and the 2 to get: 19-2x-3 = 2x+6+x Combine like terms: 16-2x=3x+6 Subtract 6 from both sides: 10-2x=3x Add 2x to both sides: 10=5x Divide by 5: 2=x